OpenStudy (anonymous):

A cylinder has a volume of 175pi cubic units and a height of 7 units. The diameter of the cylinder is

OpenStudy (goldphenoix):

Do you know the formula for finding the volume of a cylinder?

OpenStudy (anonymous):

\[V = \pi \left( \frac{ d^2 }{ 4 } \right)h\]

OpenStudy (anonymous):

I guess not, I thought it was V=Bh

OpenStudy (goldphenoix):

The formula for finding the volume of a cylinder is: \[\large \large V= hr^2\pi\]

OpenStudy (goldphenoix):

Now plug it in. \[\large \large 175 = 7r^2\pi\] What do we do first? Well, you would want to save pi for later, so that you can have an accurate answer later. So you want to get rid of 7. How can we get rid of 7?

OpenStudy (anonymous):

divide 7?

OpenStudy (anonymous):

from both sides of the equation?

OpenStudy (goldphenoix):

Good. 175 / 7 = ?

OpenStudy (anonymous):

25

OpenStudy (goldphenoix):

So now your equation looks like: \[\large \large 25 = r^2\pi\] What do you think you do next?

OpenStudy (anonymous):

square root. So the end equation would look like 5=pi?

OpenStudy (goldphenoix):

No. 5 = r(pi) Hmm. Then again, I feel like I did something wrong.

OpenStudy (goldphenoix):

Is it a multiple choice question?

OpenStudy (anonymous):

yes

OpenStudy (goldphenoix):

Can you give out the multiple choices.

OpenStudy (anonymous):

actually i cant anymore. its part of an online class and i only have a set number of attempts and it logged me out so that counted as an attempt.

OpenStudy (goldphenoix):

Oh, my bad. =/